A short quiz has two true-false questions and one multiple-choice question with four choices. A student guesses at each question. Assuming the choices are all equally likely, what is the probability that the student gets all three correct.
1/32
1/3
1/8
1/16
Please walk me thru this problem I am trying to take this online stats class and am failing miserably.
Thank you,
Kim

Answer by stanbon(75887)

A short quiz has two true-false questions and one multiple-choice question with four choices. A student guesses at each question. Assuming the choices are all equally likely, what is the probability that the student gets all three correct.
-----
The events are independent, so P(A and B) = P(A)*P(B)
---
P(2 correct on true-false) = (1/2)(1/2) = 1/4
P(1 correct mult-choice) = 1/4
----
P(this and that) = (1/4)(1/4) = 1/16
============================================
Cheers,
Stan H.
===================================
1/32
1/3
1/8
1/16